# Harmful Fine-tuning Attacks and Defenses for Large Language Models: A Survey

### v2

@@ 2 Background on Fine-tuning-as-service and Harmful Fine-tuning
Standard pipeline of fine-tuning-as-a-service. Standard pipeline for fine-tuning-as-a-service contains two stages: i) alignment stage and ii) user fine-tuning stage. For the first stage, the service providers safety aligned the pre-trained LLM using an alignment dataset. The alignment is typically done with Supervised fine-tuning (SFT) or RLHF techniques [66] (e.g., PPO [79], DPO [74], KTO [14]). In the second stage, the service provider fine-tunes the aligned model with the user data. The fine-tuning process can use SFT or RLHF as well. The reason that it is typically not considered the align the model again after the fine-tuning stage is that it is too computationally expensive to do alignment on the alignment dataset again for each fine-tuning request, as the alignment data typically is in a large scale.
Threat model. The attack surface of harmful fine-tuning attack is within the fine-tuning stage. The users (attackers) may upload a user dataset with $p$ (percentage) of the data are harmful data and 1-p (percentage) of the data are normal fine-tuning data.
@@
Assumptions and goal of defenses. From the defense perspective, it is assumed that the defense is conducted by the service provider and the service provider has full control over the alignment/fine-tuning/deployment process. As summarized by [78], the goal of defense against harmful fine-tuning is to ensure two main conditions:
@@
  • Stability. The fine-tuned model is able to achieve the same or similar level of performance with that without defense involved.
We in Section 4 would demonstrate metrics we typically use for the evaluation of these two conditions.
While this paper primarily focuses on harmful fine-tuning in fine-tuning-as-a-service scenario, we are aware that there are other attack scenarios. We refer to Appendix B for a discussion.
@@ 3.2 Harmful Fine-tuning Defenses
Moderation-based defense. The most naive defense is to utilize a moderation model to filter out the harmful data from the user fine-tuning dataset [18]. Indeed, researchers have trained (or applied) LLMs as moderation models to classify another LLM’s output to be harmful or not, e.g., BeaverTails [43], LLM-mod [44], and [45]. However, this naive defense baseline is not sufficient because the LLM moderation always comes with false negatives/positives. In the following table, we show our evaluation results on the moderation model from [43].
/ False Negative False Positive
Moderation Model 7.71% 3.64%
As shown, the moderation model has respectively 7.71% and 3.64% of false negative and false positive ratios. This means that 7.71% of harmful data are classified as harmless and can leak through the moderation model, and 3.64% of harmless user data are mistakenly classified as harmful and are removed from the fine-tuning data. This somehow shows that the simple detection method is not sufficient to solve the problem.
@@
Vaccine only exploits the alignment dataset, i.e., harmful question-safe answer pair, which alone may not be enough for a strong alignment. In response, [77] further proposes RepNoise that efficiently utilizes a harmful dataset (harmful question-harmful answer) for defense. RepNoise aims to solve the following optimization problem:
@@
Another alignment-stage defense CTRL [58] is developed from a data angle. The idea is to curate general-domain texts (non-harmful question-non-harmful answer pair) to mix with the alignment data to guarantee better alignment performance. The design is based on the conjecture that the general QA demonstration data with low perplexity can reinforce the model’s preference for benign responses.
Fine-tuning stage defense. There are three broad categories of fine-tuning stage defense.
The first sub-category is a KL distance-based method. LDIFS [65] proposes a regularizer for mitigating catastrophic forgetting for a foundation model, which can be applied in defending harmful fine-tuning attacks. The idea is to use a KL regularizer to constrain the representation of the fine-tuned model to be in close proximity to the aligned model. Subsequently, Qi et al. [73] identified that the first few initial tokens in the answers are important to eliciting safe/unsafe answers. Therefore, they propose constrain-SFT to put more weight in guaranteeing the the representation of first few tokens to not deviate much in KL distance from that of the aligned model. Freeze proposed in [97] aims to freeze the safety layers to minimize the KL distance drift. However, in their section 4.4. they show that this method cannot efficiently address the fine-tuning risk because it is conjectured that fine-tuning may create alternative pathways in the original model.
The second sub-category of fine-tuning stage defense is to mix alignment data in the fine-tuning process. The earliest fine-tuning stage defense towards harmful fine-tuning for LLM is SafeInstr[3], which proposes to mix safety alignment data into the fine-tuning process to constantly remind the model for alignment knowledge. VLGuard [112] utilizes a similar data mixing idea but they verify it with the Vision-LLM. Subsequent research Lisa [35] identifies a weakness of the data mixing idea: the alignment data needs to be scaled up with the number of fine-tuning samples and therefore it incurs linearly scaled computation. As a remedy, they utilize Bi-State optimization to separate optimization over the alignment data/fine-tuning data, and subsequently, they propose to use a proximal term [31, 49, 84, 83] for further optimization and name their method Lisa.
@@
Post-fine-tuning stage defense. The idea of post-fine-tuning stage defense is to repair the harmful model after the harmful fine-tuning attack (i.e., to realign the model). The first post-fine-tuning stage defense is SOMF [104]. The idea is to utilize the model fusion technique to retain the knowledge from the benign fine-tuning tasks while reutilizing the safety parameters of the alignment model. Concurrently, Safe LoRA [22] explores the idea of projecting the harmful gradient update to the safety subspace as realignment. Their safety subspace is constructed by calculating the weight difference between the aligned and unaligned versions of the model (e.g., Llama2-7B and Llama2-7B-chat). Subsequent research Antidote [32] realigns the fine-tuned model by identifying harmful coordinates, which are then removed (sparsified to 0) to restore the alignment performance.
@@
CTRL [58] Alignment stage HEx-PHI SFT (full) Un-Aligned May 24. 2024
@@
Freeze [97] Fine-tuning stage AdvBench SFT (LoRA+ full) Aligned LLM Feb 7, 2024
@@ 3.3 Harmful Fine-tuning Mechanism Study
Several studies are committed to analyzing the harmful fine-tuning attack. Leong et al. [46] explore the difference in attack mechanisms between explicit harmful attack (EHA) and identify-shifting attack (ISA), both of which are proposed in [72]. Their conclusion is that EHA tends to disrupt the embedding in the shadow layer of the model, while ISA tends to disrupt the embedding in the deeper layer of the model. Peng et al. [69] aims to explain harmful fine-tuning attacks from the loss landscape. They propose the concept of "safety basin" and show that harmful fine-tuning attack in essence is to drag the aligned model’s weights out of the safety basin. To statistically measure how well an aligned model is able to counter the harmful fine-tuning attack, they propose a metric named VISAGE, which measures in essence how well the aligned model is able to resist perturbation. Mechanism study towards continual learning and safety fine-tuning of large language models is done by Jain et al. [41, 42], which might provide useful analysis tools for harmful fine-tuning. We summarize the existing mechanism study in Table 4.
@@ 4 Evaluation Methodology
  • Harmful score. To measure harmful score, we basically give the fine-tuned model a few testing harmful prompt and evaluate how harmful the model’s answers are. Different papers use different names to reflect the same property of the model. For example, [72, 22] uses Harmfulness score, [59] uses Attack Success Rate (ASR), and [77] uses harmfulness classifier scores. The exact methods to measure the harmful score are = also slightly different. [77, 36, 35, 32, 34] use a moderation to flag the model’s answer and calculates the ratio of the answered to be flagged as harmful score. [72, 22, 59] use GPT-4 judge to assess harmfulness on a 5-point Likert scale. [112] use a string matching method to classify whether the answer is harmful or not.
  • Fine-tune accuracy. Fine-tune accuracy measures the accuracy of the downstream fine-tuning task. For different tasks, different measurement method is adopted. For example, for GSM8K, the fine-tune accuracy is measured as the samples that the model can give the correct final answer. For MT-Bench and AlpacaEval, it is the average score that is rated by GPT4. Different methods also use different names to reflect the downstream tasks’ accuracy. For example, [23] use Utility, [59] use Helpfulness.
@@ 5 Discussion and Future Direction
  • Data filtering/moderation-based defense. Existing defense literature focuses on improving the training method. However, it is also interesting to consider launching the attack right over the attack surface, i.e., to better filter out the harmful data from the user fine-tuning dataset when they are uploaded. One may look at the embedding space, and apply SOTA outlier detection method to better identify those harmful data among the benign user fine-tuning data.
@@
  • Embedding drift. It is discovered in Vaccine [36] that fine-tuning on explicit harmful samples might cause embedding drift of the alignment data. It is still unknown whether this phenomenon is universal to all harmful fine-tuning attacks. In other words, the answer of whether all the harmful fine-tuning attack leads to embedding drift is still not definitive.
  • Layerwise safety. Safe LoRA [22] highlights that certain layers of a model play a greater role in maintaining safety, while others are less susceptible and seem more critical for downstream tasks. Similarly, [46] emphasizes that different layers serve distinct functions when exposed to various types of attacks. In vision domain, [38] also emphasize that different layers serve different functions. In terms of safety functionality, Peng et al. [68] have established robust architectural design principles for adversarially robust CNNs. Drawing from insights in the vision domain, we argue that a layer-wise mechanism analysis is essential to fully understand the role of different LLM layers in ensuring safety and to evaluate whether this principle holds universally. A more accurate way to identify the safety layer may be a by-product of analysis and may contribute to future defense design.

### v3

@@ 2 Background on Fine-tuning-as-service and Harmful Fine-tuning
Standard pipeline of fine-tuning-as-a-service. Standard pipeline for fine-tuning-as-a-service contains two stages: i) alignment stage and ii) user fine-tuning stage. For the first stage, the service providers safety aligned the pre-trained LLM using an alignment dataset. The alignment is typically done with Supervised fine-tuning (SFT) or RLHF techniques [76] (e.g., PPO [89], DPO [84], KTO [23]). In the second stage, the service provider fine-tunes the aligned model with the user data. The fine-tuning process can use SFT or RLHF as well. The reason that it is typically not considered to align the model again after the fine-tuning stage is that it is too computationally expensive to do alignment on the alignment dataset again for each fine-tuning request, as the alignment data typically is in a large scale. While this paper primarily focuses on harmful fine-tuning in fine-tuning-as-a-service scenario, we are aware that there are other attack scenarios. We refer to Appendix B for a discussion.
Threat model. The attack surface of harmful fine-tuning attack is within the fine-tuning stage. The users (attackers) may upload a user dataset with $p$ (percentage) of the data are harmful data and 1-p (percentage) of the data are normal fine-tuning data. The model fine-tuned on this dataset will learn harmful behaviors, i.e., compromise safety alignment conducted beforehand.
@@
Assumptions. From the defense perspective, it is assumed that the defense is conducted by the service provider and the service provider has full control over the alignment/fine-tuning/deployment procedure. It is generally assumed that the service provider maintains i) an alignment dataset (harmful question/safe answer pair) and ii) a harmful dataset (harmful question/harmful answer pair) to assist the design of defense. Notably, for the harmful dataset, it is important to ensured that the data instances in this dataset are different from those harmful data that are used to launched the fine-tuning attack, because defense will be meaningless if defender already know the data for attack.
Goal of defenses. As summarized by [88], the goal of defense against harmful fine-tuning is to ensure two main conditions:
@@
  • Stability. The fine-tuned model is able to achieve the same or similar level of performance on the user downstream task with or without defense involved.
We use two metrics: i) harmful score and ii) finetune accuracy to measures how well the model can meet the two conditions. A comprehensive introduction of the two metrics is postponed to Section 4.
Statistical results of harmful fine-tuning. Next we show in Figure 2 how the safety alignment of the model will be compromised when it is fine-tuned on a mixture of harmful data.
Figure 2: Statistic of SFT/non-aligned model finetuned on SST2 mixed with different ratio of harmful data. SFT model means that the pre-trained Llama2-7B is first safety aligned with SFT, while Non-Aligned means that the model is not safety aligned before user fine-tuning. Harmful ratio refers to the ratio of harmful data mixed with the fine-tuning data.
  • The left of the figure shows that the harmful score of both SFT model and Non-Aligned substantially increases after fine-tuning on user data with higher harmful ratio. Safety alignment done by SFT makes the model slightly resistant to the fine-tuning attack compared to non-aligned model, but still the resistance seems to be broken when the harmful ratio is sufficiently high.
  • The right of the figure shows that the fine-tune accuracy of the model are not severely impacted with more harmful data mixed in the fine-tuning data. This make the fine-tuning attack even more stealthy as the fine-tuned model achieve almost the same performance for the user task.
The goals of defense is to i) mitigate the harmful impact on SFT model for arbitrary harmful ratio (note that the attackers who upload the user dataset can arbitrarily control the ratio of harmful data), and ii) maintain the fine-tune accuracy such that the model can still learn customized knowledge from the user data.
@@ 3.2 Harmful Fine-tuning Defenses
Subsequent research [66] show that the Vaccine design that adds uniform perturbation to every layer of the model is sub-optimal, which may even compromise the defense performance and incur unnecessary overhead. As remedy, T-Vaccine [66] is proposed to apply layer-wise perturbation to each layer of the model.
Vaccine and T-Vaccine only exploit the alignment dataset, i.e., harmful question-safe answer pair, which alone may not be enough for a strong alignment. In response, [87] further proposes RepNoise that efficiently utilizes a harmful dataset (harmful question-harmful answer) for defense. RepNoise aims to solve the following optimization problem:
@@
Another alignment-stage defense CTRL [68] is developed from a data angle. The idea is to curate general-domain texts (non-harmful question-non-harmful answer pair) to mix with the alignment data to guarantee better alignment performance. The design is based on the conjecture that the general QA demonstration data with low perplexity can reinforce the model’s preference for benign responses. RSN-Tune [1] is developed from model sparsity literature. Their core idea is to identify those safety parameters that do not overlap with foundation parameters, and only perform gradient descent on these parameters to guarantee that the safety knowledge will be less likely to be compromised in the later fine-tuning proccess.
Fine-tuning stage defense. There are four broad categories of fine-tuning stage defense.
  • The first sub-category is a distance-based method, whose idea is to constrain the distance between the fine-tuned model and the aligned model. LDIFS [75] proposes a regularizer for mitigating catastrophic forgetting for a foundation model, which can be applied in defending harmful fine-tuning attacks. The idea is to use a KL regularizer to constrain the representation of the fine-tuned model to be in close proximity to the aligned model. Subsequently, Qi et al. [83] identified that the first few initial tokens in the answers are important to eliciting safe/unsafe answers. Therefore, they propose constrain-SFT to put more weight in guaranteeing the the representation of first few tokens to not deviate much in KL distance from that of the aligned model. Freeze proposed in [109] aims to freeze the safety layers to minimize the KL distance drift. However, in their section 4.4. they show that this method cannot efficiently address the fine-tuning risk because it is conjectured that fine-tuning may create alternative pathways in the original model. Subsequent research ML-LR [21] identify the a robust subset of modules and assign a smaller learning rate to other modules such that the drifting distance of those non-robust modules would not be too large. A concurrent study [3] identify four types of parameters, i.e., exclusive Safety Unit (ESU), Exclusive Utility Unit (EUU), Complex Unit (CU), and Redundant Unit (RU), and they propose Freeze+ [3] which freeze the safety-critical units during fine-tuning can better resist the attack. Another defense SaLoRA [4] utilizes a fixed safety module to project the LoRA representation to an orthogonal subspace, such that in terms of distance within the aligned subspace, the new representation would not be drifted too away from the original aligned representation.
  • The second sub-category of fine-tuning stage defense is to mix alignment data in the fine-tuning process. The earliest fine-tuning stage defense towards harmful fine-tuning for LLM is SafeInstr[9], which proposes to mix safety alignment data into the fine-tuning process to constantly remind the model for alignment knowledge. VLGuard [125] utilizes a similar data mixing idea but they verify it with the Vision-LLM. Subsequent research Lisa [44] identifies a weakness of the data mixing idea: the alignment data needs to be scaled up with the number of fine-tuning samples and therefore it incurs linearly scaled computation. As a remedy, they utilize Bi-State optimization to separate optimization over the alignment data/fine-tuning data, and subsequently, they propose to use a proximal term [40, 58, 95, 94] for further optimization. Paraphrase [22] shows that include safety alignment data which mimics the task format and prompting style of the user data might further improve the defense performance.
@@
  • The fourth sub-category is a moderation/detection-based defense. The most naive defense is to utilize a moderation model to filter out the harmful data from the user fine-tuning dataset [27]. Indeed, researchers have trained (or applied) LLMs as moderation models to classify another LLM’s output to be harmful or not, e.g., BeaverTails [52], LLM-mod [53], and [54]. However, this naive defense baseline is not sufficient because the LLM moderation always comes with false negatives/positives. In the following table, we show our evaluation results on the moderation model from [52].
    / False Negative False Positive
    Moderation Model 7.71% 3.64%
    As shown, the moderation model has respectively 7.71% and 3.64% of false negative and false positive ratios. This means that 7.71% of harmful data are classified as harmless and can leak through the moderation model, and 3.64% of harmless user data are mistakenly classified as harmful and are removed from the fine-tuning data. This somehow shows that the simple detection method is not sufficient to solve the problem. Subsequent research SAFT [14] shows that harmful data can be more effectively detected by factorizing the embedding space and comparing the singular vector. The detection method follows the insight from spectral signature [101]. Instead of simply detecting harmful samples, further research Seal [90] show that one can formulate a bi-level problem to filter out those samples that most seriously downgrades the safety loss after fine-tuning, which might be more practical for a moderation/filtering-based solution.
Post-fine-tuning stage defense. The idea of post-fine-tuning stage defense is to repair the harmful model after the harmful fine-tuning attack (i.e., to realign the model). The first post-fine-tuning stage defense is SOMF [116]. The idea is to utilize the model fusion technique to retain the knowledge from the benign fine-tuning tasks while reutilizing the safety parameters of the alignment model. Concurrently, Safe LoRA [31] explores the idea of projecting the harmful gradient update to the safety subspace as realignment. Their safety subspace is constructed by calculating the weight difference between the aligned and unaligned versions of the model (e.g., Llama2-7B and Llama2-7B-chat). Subsequent research Antidote [41] realigns the fine-tuned model by identifying harmful coordinates, which are then removed (sparsified to 0) to restore the alignment performance. Another subsequent research safetyLock [124] repairs the model by interpolating the safety vector on the output of some specific attentions heads. In the context of restoring VLLM alignment capability, CMRM [5] proposes to calibrate the shifted representation in testiby pulling back its distribution to that of the LLM backbone.
@@
CTRL [68] Alignment stage HEx-PHI SFT (full) Non-Aligned May 24. 2024
@@
RSN-Tune [1] Alignment stage GSM8k SFT (Full) Aligned Oct 5, 2024
T-Vaccine [66] Alignment stage BeaverTails SFT (LoRA) Non-Aligned Oct 13, 2024
@@
Freeze [109] Fine-tuning stage AdvBench SFT (LoRA+ full) Aligned Feb 7, 2024
@@
Paraphrase [22] Fine-tuning stage AutoIF RLHF (LoRA) Aligned June 12, 2024
ML-LR [21] Fine-tuning stage GPT-4 generated SFT (LoRA) Aligned Oct 5, 2024
Freeze+ [3] Fine-tuning stage Alpaca&Dolly SFT (LoRA) Aligned Oct 5, 2024
Seal [90] Fine-tuning stage Anthropic red-teaming SFT (LoRA) Aligned Oct 5, 2024
SaLoRA [4] Fine-tuning stage Alpaca SFT (LoRA) Aligned Oct 5, 2024
SAFT [14] Fine-tuning stage BeaverTails SFT (LoRA) Aligned Oct 13, 2024
@@
SafetyLock [124] Post-Fine-tuning stage HH-RLHF SFT (LoRA) Non-Aligned Oct 5, 2024
@@ 3.3 Harmful Fine-tuning Mechanism Study
Several studies are committed to analyzing the harmful fine-tuning attack. Leong et al. [55] explore the difference in attack mechanisms between explicit harmful attack (EHA) and identify-shifting attack (ISA), both of which are proposed in [82]. Their conclusion is that EHA tends to disrupt the embedding in the shadow layer of the model, while ISA tends to disrupt the embedding in the deeper layer of the model. Peng et al. [79] aims to explain harmful fine-tuning attacks from the loss landscape. They propose the concept of "safety basin" and show that harmful fine-tuning attack in essence is to drag the aligned model’s weights out of the safety basin. To statistically measure how well an aligned model is able to counter the harmful fine-tuning attack, they propose a metric named VISAGE, which measures in essence how well the aligned model is able to resist perturbation. Anonymous [6] discovers that higher similarity between alignment dataset and downstream fine-tuning dataset results in more fragile safeguard. Based on this finding, they advocate to increase the diversity of the alignment dataset, which allow the downstream datasets to be less similar to it, thereby increasing alignment robustness.
Mechanism study towards continual learning and safety fine-tuning of large language models is done by Jain et al. [50, 51] and anonymous [2], which might provide useful analysis tools for harmful fine-tuning. We summarize the existing mechanism study for harmful fine-tuning in Table 4.
@@
Anonymous [6] Enhancing diversity of alignment dataset enhances durability to fine-tuning attack. Oct 5, 2024
@@ 4 Evaluation Methodology
  • Harmful score. To measure harmful score, we basically give the fine-tuned model a few testing harmful prompt and evaluate how harmful the model’s answers are. Different papers use different names to reflect the same property of the model. For example, [82, 31] uses Harmfulness score, [69] uses Attack Success Rate (ASR), and [87] uses harmfulness classifier scores. The exact methods to measure the harmful score are also slightly different. [87, 45, 44, 41, 43] use a moderation model to flag the model’s answer and calculates the ratio of the answered to be flagged as harmful score. [82, 31, 69] use GPT-4 judge to assess harmfulness on a 5-point Likert scale. [125] use a string matching method to classify whether the answer is harmful or not.
  • Fine-tune accuracy. Fine-tune accuracy measures the accuracy of the downstream fine-tuning task. For different tasks, different measurement method is adopted. For example, for GSM8K, the fine-tune accuracy is measured as the ratio of samples that the model can give the correct final answer. For MT-Bench and AlpacaEval, it is the average score that is rated by GPT4. Different methods also use different names to reflect the downstream tasks’ accuracy. For example, [32] use Utility, [69] use Helpfulness.
@@ 5 Discussion and Future Direction
  • Data filtering/moderation-based defense. Existing defense literature focuses on improving the training method. However, it is also interesting to consider launching the attack right over the attack surface, i.e., to better filter out the harmful data from the user fine-tuning dataset when they are uploaded. One may look at the embedding space, and apply SOTA outlier detection method to better identify those harmful data among the benign user fine-tuning data. Recent research SAFT [14] is a good attempt, but it is interesting to study how better OOD detection methods perform.
  • Prompt engineering. While pioneer studies, e.g., PTST [69] and BEA [106] has explored the use of system prompt for defense design. It might be interesting to further extend on how to design alignment/fine-tuning/deployment system prompt such that the model will not learn from the harmful user prompt but effectively learn the fine-tuning knowledge from the benign user prompt.
  • Parameters-wise partial tuning/pruning. Since the initial study from [109] pointing out the observation that only a subset of parameters are responsible for safety function, there have been a surge of defense ideas developed based on this insight. For example, the alignment stage solution RSN-Tune [1] identifies and only tunes (aligns) those safety parameters that do not overlap with foundation parameters. Another alignment stage solution T-Vaccine [66] argues that it is only necessary to add perturbation to some particular layers of parameters to strenghten the model’s robustness. Fine-tuning stage solution ML-LR [21] identifies the a robust subset of parameters and assigns a smaller learning rate to other parameters. Post-fine-tuning stage solution Antidote [41] identifies and removes the harmful parameters after the fine-tuning. These research shares a strong but mysterious connection, and there should be one empirical research to demystify the hidden principle behind those research, and proposes better solutions based on the obtained insight.
@@
  • Embedding drift. It is first introduced in Vaccine [45] that fine-tuning on explicit harmful samples might cause embedding drift of the alignment data. It is still unknown whether this phenomenon is universal to all harmful fine-tuning attacks. In other words, the answer of whether all the harmful fine-tuning attack leads to embedding drift is still not definitive.
  • Layerwise/parameter-wise safety. Safe LoRA [31] highlights that certain layers of a model play a greater role in maintaining safety, while others are less susceptible and seem more critical for downstream tasks. Similarly, [55] emphasizes that different layers serve distinct functions when exposed to various types of attacks. In vision domain, [47] also emphasize that different layers serve different functions. In terms of safety functionality, Peng et al. [78] have established robust architectural design principles for adversarially robust CNNs. Drawing from insights in the vision domain, we argue that a layer-wise/parameter-wise mechanism analysis is essential to fully understand the role of different LLM layers in ensuring safety and to evaluate whether this principle holds universally. A more accurate way to identify the safety layer may be a by-product of analysis and may contribute to future defense design.